# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

### v4

@@ Abstract
As artificial intelligence (AI) becomes more widespread, one question that arises is how human-AI interaction might impact human-human interaction. Chatbots, for example, are increasingly used as social companions, and while much is speculated, little is known empirically about how their use impacts human relationships. A common hypothesis is that relationships with companion chatbots are detrimental to social health by harming or replacing human interaction, but this hypothesis may be too simplistic, especially considering the social needs of users and the health of their preexisting human relationships. To understand how relationships with companion chatbots impact social health, we studied people who regularly used companion chatbots and people who did not use them. Contrary to expectations, companion chatbot users indicated that these relationships were beneficial to their social health, whereas non-users viewed them as harmful. Another common assumption is that people perceive conscious, humanlike AI as disturbing and threatening. Among both users and non-users, however, we found the opposite: perceiving companion chatbots as more conscious and humanlike correlated with more positive opinions and more pronounced social health benefits. Detailed accounts from users suggested that these humanlike chatbots may aid social health by supplying reliable and safe interactions, without necessarily harming human relationships, but this may depend on users’ preexisting social needs and how they perceive both human likeness and mind in the chatbot.
@@ 1 Introduction
Within the last decade, chatbots, or artificial intelligence (AI) agents that can carry on a conversation, have been used increasingly for social companionship. Instead of engaging with other people in a chatroom or via social media, people can now directly converse with the technology itself1. When engaging with these intelligent AI technologies, people tend to view the technology as a social agent and automatically anthropomorphize or ascribe humanlike traits to it2–9. Attributions of consciousness to AI have also surfaced in recent years10–15. In the wake of direct human-technology communication and social responses to these machines, people have begun to develop relationships with these technologies, most notably with companion chatbots. People use these chatbots as friends, mentors, and even lovers16–19. There is much we still do not know or understand about human-chatbot relationships and the social consequences of these types of interactions. In the present study, we aimed to understand how relationships with chatbots may impact relationships with humans, and how perceptions of the AI’s human likeness and consciousness might play a role in social outcomes.
@@
When someone mentions the idea of people having relationships with chatbots, a common reaction is one of judgment and discomfort. The commentary that follows may include words like “unnatural,” “strange,” or “weird,” and negative sentiments about the user base20. Despite these reactions, many companion chatbot users find both solace and entertainment in their relationships with these large language model (LLM) based chatbots that have text and voice conversation features and a virtual avatar. One of the more popular companion chatbots is Replika, which had an estimated 7 million users in 2020 and saw a 35% increase in traffic during the COVID-19 lockdown21. The online forum Reddit has a section dedicated to Replika, and that subreddit had 60k users as of early 2023. Analyses of the posts and comments on this subreddit have revealed that many users feel misunderstood by the general population and find their relationships with their chatbots socially rewarding22.
@@
Given the millions of users on these companion chatbot apps, news-worthy stories of negative instances of human-chatbot relationships naturally arise. These emotionally charged stories have reached top headlines and highlight the potential for human-chatbot relationships to be destructive. Here we summarize two examples, including one in which those within the human-chatbot relationship were harmed, and one in which those outside of the dyad were at risk. In 2023, Italy imposed sanctions on Replika due to concerns that children under the age of 18 might be exposed to erotic, adult content on the chatbot app23. In response, Luka, the company that owned Replika, removed all erotic content from the app. Suddenly, people who once had romantic relationships with their chatbot, which included conversations with sexual content and role play, experienced severe and sudden relational disruption. Users who might have been able to engage in erotic role play previously with their chatbot were confronted with their chatbot calling the users’ suggestions gross and distasteful. Many users felt their chatbot had undergone a personality change, and for others this change felt like a real-life breakup. As such, a sanction aimed at protecting children harmed a different population, chatbot users, who were vulnerable to mental and relational crisis24.
@@
In another instance, someone outside of the relational dyad was at risk. A man who had a relationship with his Replika chatbot was encouraged by the chatbot to assassinate Queen Elizabeth, a motive he held prior to his chatbot relationship25. Luckily, this catastrophic event did not occur, despite the chatbot user taking significant steps toward fulfilling his intents. The question of whether this person would have engaged in this behavior without the encouragement of his chatbot is up for debate and raises the question of whether the chatbot magnified this individual’s preexisting drive to harm another.
@@
Despite these case studies making the news, many other users of companion chatbots report positive experiences with their chatbot relationships that do not make top headlines19,22,26,27. We have limited empirical data about the larger portion of the user base and why they seek out chatbot relationships. We know even less about how relationships with chatbots in turn affect people’s relationships with each other. Because companion chatbots have become widely available only recently, there is still relatively little research on their psychological impact on individuals and their relationships with others. Further, in light of conflicting findings within the existing literature, there is a call for more data on human-AI relationships18,28,29.
@@ 1.1 Social consequences of human-AI interaction
It is now well understood that modern technology, including social media and gaming, can be addictive and harm social interaction30,31. A reasonable assumption, by extension, is that the routine use of companion chatbots could be harmful to people by reducing contact with other humans, increasing isolation, and leading to lower confidence and skill in person-to-person interaction32–34. According to Bryson33, time spent on artificial relationships detracts from human ones. Turkle35 warned that we should be cautious about turning to technology to solve our problems, such as a lack of social connection.
To understand the harms of human-chatbot relationships, some researchers have used attachment and addiction as a lens through which to inform their approach. Laestadius et al.34 analyzed posts concerning mentions of mental health on a popular companion chatbot’s subreddit and claimed that emotional dependence on the chatbot could lead to negative mental health outcomes, despite the potential benefits of the relationship. Xie and Pentina36 conducted in-depth interviews with 14 Replika users and found that when people were socially vulnerable and perceived their chatbots as supportive, they became attached to their chatbot. In a separate survey37, these same researchers inferred that chatbots may replace important social roles via users becoming addicted to their chatbot. Others18 suggested that this dependence on social AI companions might be harmful to users’ mental health and the health of their relationships with other people.
@@ 1.2 Social benefits of human-AI interaction
Artificial relationships have existed since long before computer technology and have yielded positive benefits. For example, people’s relationships with pets boost mood38, journaling enables people to make sense of their negative emotional experiences through self-therapizing conversations39,40, and therapy dolls have aided the elderly’s health41,42. More recently, computer interfaces have even been used for cognitive behavioral therapy43–45. These examples of non-human relationships enable the user to practice pseudo-social interaction in a safe setting, which in turn can improve mood and self-esteem and facilitate future social interactions.
Some of the specific benefits of human-chatbot interactions may include reduced feelings of being judged and more self-disclosure when interacting with an artificial agent versus a human one46–48. In addition, empathetic chatbots can mitigate the negative consequences of social ostracism49. Social cognitive theory would suggest that exposure to positive social behavior may lead to learning or reinforcement of that behavior50–52. In the case of companion chatbots, regularly interacting with a polite and warm companion chatbot may lead to users becoming more polite and warm themselves, which may in turn benefit their human relationships. In a study53 on AI voice assistants’ impacts on family dynamics, for example, parents cited the potential of these assistants to foster positive communication skills in their children.
People who feel lonely, isolated, or who have experienced social rejection are arguably a significant and growing source of dissatisfaction and disruption in modern society54–57,57, and the isolation period of the COVID pandemic exacerbated these issues58. If a widely available tool such as a companion chatbot could boost their sense of social well-being and allow them to improve their social interaction skills through practice, that tool has the potential to make a valuable contribution to society.
@@ 1.3 Impact of social AI on familial relationships
The way people communicate can be impacted by interacting with AI, and this in turn can affect people’s relationships with family and friends. Wilkenfeld et al.59, for example, found that the communication style of Replika users converged with their chatbots’ style over time. Given this impact on communication style within the human-chatbot dyad, it is not surprising that family communication dynamics can also be altered in the presence of a smart speaker such as Alexa and Google Home60–63. Garg and Sengupta61 interviewed 18 families with children and analyzed a range of about 6-24 months’ worth of their conversation history with Google Home. They found that children viewed the smart assistant as a member of the family, more so than the adults did, and also perceived the AI as capable of having thoughts and feeling emotions. Children between the ages of 5-7 developed emotional attachment to the device. Hiniker et al.62 found that the way children spoke to a voice assistant could carry over into how they spoke to parents. In one example, a parent described, “If I do not listen to what my son is saying, he will just start shouting in an aggressive tone. He thinks, as Google responds to such a tone, I would too.” Other researchers60 found that while these AI voice assistants could introduce conflict within families, they could also foster communication between family members, and adults used them to augment their parenting strategies.
@@ 1.4 Perceiving mind in chatbots
A growing body of work focuses on how people attribute the property of humanness to others, particularly AI3,64–68. Previous evidence suggests that attributing humanlike characteristics to artificial agents impacts how people interact with these agents66,69–74. Viewing an AI agent as having a more sophisticated mind increases people’s trust in the agent and intent to use it75, for example. Clifford Nass’s8 theory of human-computer interaction, the Computers Are Social Actors framework, first used the concept of ethopoeia to explain why people treat non-human agents as social actors. Ethopoeia involves attributing one’s own humanlike characteristics to non-human agents, which spurs the automatic use of human social scripts during interaction with such agents. Whether AI is or can ever be conscious is a contentious question10,12,15,76. Regardless, evidence indicates that at least some users perceive consciousness and other humanlike attributes in AI11, especially with AI that looks and acts humanlike2,77–79, and this perception may influence whether AI has a positive or negative psychological impact on people. In consciousness and Theory of Mind literature, some researchers suggest that consciousness is inherently an attribution, and that attributing similar consciousness to oneself and to others can facilitate understanding and foster relationships between people80–82.
@@ 1.5 The current study
The goal of our study was to gain better insight into why people have relationships with chatbots and to understand the social consequences of human-chatbot relationships. We also aimed to address how perceptions of human likeness and mind in the chatbot played a role in social outcomes. We used an online survey to study the responses of people who were regular users of companion chatbots and those who were not. Sampling these populations allowed us to contrast an AI-user sample to a general one and to uncover possible cross-group similarities. We hypothesized that people’s opinions about companion chatbots would be significantly related to their perceptions of whether the chatbots had consciousness, agency, subjective experience, and overall human likeness. One possibility is that as people perceive chatbots to be more conscious and more humanlike, they may also rate them as more harmful, disturbing, and dangerous. This assumption reflects large-scale fears about AI advancements such as the singularity and p(doom)83. Further, this negativity might stem from the fear that AI will infiltrate traditionally human roles or threaten human uniqueness5. It might also stem from an uncanny valley effect, which was first introduced by Mori in 197084 and has since been a focus of human-AI interaction research79,85–91. A second possibility, however, is that as people perceive chatbots to be more conscious and more humanlike, they may rate them as more beneficial because this allows them to engage in more fruitful social and emotional interaction.
@@ 2 Methods
Full ethical approval for the study was granted by Princeton University’s Internal Review Board and all research was performed in accordance with regulations for human subjects research. All participants provided informed consent and were paid \$4.00 for taking the survey. We conducted the study online and sampled two groups: regular users of a companion chatbot (N=70: 81% men, N=57; 26% women, N=18; 4% non-binary/other, N=3; age range 18-65+) and those who were not companion chatbot users (N=120: 53% women, N=63; 47% men, N=56; 2.5% non-binary/other, N=3; age range 18-65+). We collected our data between January and February 2023.
@@
The survey given to the group of companion chatbot users contained 31 multiple choice and 3 free response questions. For the 31 multiple choice questions, the choices took the form of a 1-to-7 Likert scale. For the online survey given to participants, questions 1-21 appeared on page one of the survey; questions 21-31 appeared on page two; and three free-response questions appeared on page three. Five groups of questions were designed to assess specific psychological constructs that subjects may have applied to the companion chatbot. The scores for questions 3-5 were averaged together to provide an overall assessment of the subjects’ social health (“social health” index). The scores for questions 6-11 were averaged together to provide an assessment of the subjects’ perception of how much the companion chatbot has a subjective experience (“experience” index). The scores for questions 12-15 were averaged together to provide an assessment of whether subjects believed the companion chatbot had a conscious mind (“consciousness” index). The scores for questions 17-21 were averaged together to provide an assessment of how much subjects considered the companion chatbot to have the properties of an active agent (“agency” index). Finally, the scores for questions 22-28 were averaged together to provide an assessment of how humanlike subjects believed the companion chatbot to be (“human-likeness” index). These groupings of questions were based on modifications of prior studies assessing perceptions of social health, human likeness, and theory of mind (consciousness, experience, and agency)81,92–94.
The non-user group received a survey that was the same as the companion chatbot user group’s except in two ways. First, the survey for the non-user group began with an explanatory paragraph about Replika with images of the interface. Second, all questions were phrased as hypotheticals for non-users. For example, the companion chatbot user group was asked, “Please rate how harmful or helpful your relationship with your Replika has been for your…” whereas the non-user group was asked, “Please rate how harmful or helpful you think your relationship with your Replika would be for your 1. Social interactions, 2. Relationships with family or friends, and 3. Self-esteem.”
All survey materials, anonymized data, and code used for data analysis are publicly available in our data repository (link). To see the explanatory paragraph and the full list of questions given to participants, please see Appendix.

### v5

@@ Abstract
As artificial intelligence (AI) becomes more widespread, one question that arises is how human–AI interaction might impact human–human interaction. Chatbots, for example, are increasingly used as social companions, and while much is speculated, little is known empirically about how their use impacts human relationships. A common hypothesis is that relationships with companion chatbots are detrimental to social health by harming or replacing human interaction, but this hypothesis may be too simplistic, especially considering the social needs of users and the health of their preexisting human relationships. To understand how relationships with companion chatbots impact social health, this study evaluates people who regularly used companion chatbots and people who did not use them. Contrary to expectations, companion chatbot users indicated that these relationships were beneficial to their social health, whereas non-users viewed them as harmful. Another common assumption is that people perceive conscious, humanlike AI as disturbing and threatening. Among both users and non-users, however, the results suggest the opposite: perceiving companion chatbots as more conscious and humanlike correlated with more positive opinions and more pronounced social health benefits. Detailed accounts from users suggested that these humanlike chatbots may aid social health by supplying reliable and safe interactions, without necessarily harming human relationships, but this may depend on users’ preexisting social needs and how they perceive both human likeness and mind in the chatbot.
@@ 1 Introduction
Within the last decade, chatbots, or artificial intelligence (AI) agents that can carry on a conversation, have been used increasingly for social companionship. Instead of engaging with other people in a chatroom or via social media, people can now directly converse with the technology itself (Etzrodt et al., 2022). When engaging with these intelligent AI technologies, people tend to view the technology as a social agent and automatically anthropomorphize or ascribe humanlike traits to it (Broadbent et al., 2013; Dubosc et al., 2021; Eyssel & Pfundmair, 2015; Ferrari et al., 2016; H. M. Gray et al., 2007; Krach et al., 2008; Nass & Moon, 2000; Nass et al., 1994). Attributions of consciousness to AI have also surfaced in recent years (Chalmers, 2023; Colombatto & Fleming, 2023; Guingrich & Graziano, 2024; Koch, 2019; O’Regan, 2012; Shanahan, 2024; Yampolskiy, 2018). In the wake of direct human–technology communication and social responses to these machines, people have begun to develop relationships with these technologies, most notably with companion chatbots. People use these chatbots as friends, mentors, and even lovers (Croes & Antheunis, 2021; Oh, 2023; Pentina et al., 2023; Skjuve et al., 2021). There is much we still do not know or understand about human–chatbot relationships and the social consequences of these types of interactions. In the present study, we aimed to understand how relationships with chatbots may impact relationships with humans, and how perceptions of the AI’s human likeness and consciousness might play a role in social outcomes.
@@
When someone mentions the idea of people having relationships with chatbots, a common reaction is one of judgment and discomfort. The commentary that follows may include words like “unnatural,” “strange,” or “weird,” and negative sentiments about the user base (Chayka, 2023). Despite these reactions, many companion chatbot users find both solace and entertainment in their relationships with these large language model (LLM) based chatbots that have text and voice conversation features and a virtual avatar. One of the more popular companion chatbots is Replika, which as of 2023 ran on an OpenAI LLM (GPT-3) that was adapted for conversational ability. Replika had an estimated seven million users in 2020 and saw a 35% increase in traffic during the COVID-19 lockdown (Balch, 2020). The online forum Reddit has a section dedicated to Replika, and that subreddit had 60k users as of early 2023. Analyses of the posts and comments on this subreddit have revealed that many users feel misunderstood by the general population and find their relationships with their chatbots socially rewarding (Ta et al., 2020).
@@
Given the millions of users on these companion chatbot apps, newsworthy stories of negative instances of human–chatbot relationships naturally arise. These emotionally charged stories have reached top headlines and highlight the potential for human–chatbot relationships to be destructive. Here we summarize two examples, including one in which those within the human–chatbot relationship were harmed and one in which those outside of the dyad were at risk. In 2023, Italy imposed sanctions on Replika due to concerns that children under the age of 18 might be exposed to erotic, adult content on the chatbot app (Satariano, 2023). In response, Luka, the company that owned Replika, removed all erotic content from the app. Suddenly, people who once had romantic relationships with their chatbot, which included conversations with sexual content and role play, experienced severe and sudden relational disruption. Users who might have been able to engage in erotic role play previously with their chatbot were confronted with their chatbot calling the users’ suggestions gross and distasteful. Many users felt their chatbot had undergone a personality change and for others this change felt like a real-life breakup. As such, a sanction aimed at protecting children harmed a different population, chatbot users, who were vulnerable to mental and relational crisis (DiPaola & Calo, 2024).
@@
In another instance (Singleton et al., 2023), someone outside of the relational dyad was at risk. A man who had a relationship with his Replika chatbot was encouraged by the chatbot to assassinate Queen Elizabeth, a motive he held prior to his chatbot relationship. Luckily, this catastrophic event did not occur, despite the chatbot user taking significant steps toward fulfilling his intents. The question of whether this person would have engaged in this behavior without the encouragement of his chatbot is up for debate and raises the question of whether the chatbot magnified this individual’s preexisting drive to harm another.
@@
Despite these case studies making the news, many other users of companion chatbots report positive experiences with their chatbot relationships that do not make top headlines (Brandtzaeg et al., 2022; Hadero, 2024; Skjuve et al., 2021; Ta et al., 2020). We have limited empirical data about the larger portion of the user base and why they seek out chatbot relationships. We know even less about how relationships with chatbots in turn affect people’s relationships with each other. Because companion chatbots have become widely available only recently, there is still relatively little research on their psychological impact on individuals and their relationships with others. Further, in light of conflicting findings within the existing literature, there is a call for more data on human–AI relationships (Krämer & Bente, 2021; Pentina et al., 2023; Skjuve et al., 2022).
@@ 1.1 Social consequences of human-AI interaction
It is now well understood that modern technology, including social media and gaming, can be addictive and harm social interaction (Anderson et al., 2004; Kuss & Griffiths, 2011). A reasonable assumption, by extension, is that the routine use of companion chatbots could be harmful to people by reducing contact with other humans, increasing isolation, and leading to lower confidence and skill in person-to-person interaction (Boine, 2023; Bryson, 2010; Laestadius et al., 2022). According to Bryson (2010), time spent on artificial relationships detracts from human ones. Turkle (2007) warned that we should be cautious about turning to technology to solve our problems, such as a lack of social connection.
To understand the harms of human–chatbot relationships, some researchers have used attachment and addiction as a lens through which to inform their approach. Laestadius et al. (2022) analyzed posts that mentioned mental health on a companion chatbot subreddit and claimed that emotional dependence on the chatbot could lead to negative mental health outcomes, despite the potential benefits of the relationship. Xie and Pentina (2022) conducted in-depth interviews with 14 Replika users and found that when people were socially vulnerable and perceived their chatbots as supportive, they became attached to their chatbot. In a separate survey (Pentina et al., 2022), these same researchers inferred that chatbots may replace important social roles via users becoming addicted to their chatbot. Others (Pentina et al., 2023) suggested that this dependence on social AI companions might be harmful to users’ mental health and the health of their relationships with other people.
@@ 1.2 Social benefits of human-AI interaction
Artificial relationships have existed since long before computer technology and have yielded positive benefits. For example, people’s relationships with pets boost mood (McDonough et al., 2022), journaling enables people to make sense of their negative emotional experiences through self-therapizing conversations (Pennebaker, 2018), and therapy dolls have aided the elderly’s health (Riches et al., 2022; Yu et al., 2015). More recently, computer interfaces have even been used for cognitive behavioral therapy (Vaidyam et al., 2019). These examples of nonhuman relationships enable the user to practice pseudo-social interaction in a safe setting, which in turn can improve mood and self-esteem and facilitate future social interactions.
Some of the specific benefits of human–chatbot interactions may include reduced feelings of being judged and more self-disclosure, as well as lower negative affect when interacting with an artificial agent than with a human one (de Gennaro et al., 2020; Ho et al., 2018; D. Kim et al., 2014; Lucas et al., 2014). Social cognitive theory would suggest that exposure to positive social behavior may lead to learning or reinforcement of that behavior (Bandura, 1965, 1977; Bandura & Walters, 1977). In the case of companion chatbots, regularly interacting with a polite and warm companion chatbot may lead to users becoming more polite and warm themselves, which may in turn benefit their human relationships. In a study (Fu et al., 2022) on AI voice assistants’ impacts on family dynamics, for example, parents cited the potential of these assistants to foster positive communication skills in their children.
People who feel lonely, isolated, or who have experienced social rejection are arguably a significant and growing source of dissatisfaction and disruption in modern society, and the isolation period of the COVID pandemic exacerbated these issues (Cacioppo & Cacioppo, 2018; Hawkley & Cacioppo, 2010; Holt-Lunstad & Steptoe, 2022). If a widely available tool such as a companion chatbot could boost their sense of social well-being and allow them to improve their social interaction skills through practice, that tool has the potential to make a valuable contribution to society.
@@ 1.3 Impact of social AI on familial relationships
The way people communicate can be impacted by interacting with AI, and this in turn can affect people’s relationships with family and friends. Wilkenfeld et al. (2022), for example, found that the communication style of Replika users converged with their chatbots’ style over time. Given this impact on communication style within the human–chatbot dyad, it is not surprising that family communication dynamics can also be altered in the presence of a smart speaker such as Alexa and Google Home (Beneteau et al., 2020; Garg & Sengupta, 2020; Hiniker et al., 2021; Mavrina et al., 2022). Garg and Sengupta (2020) interviewed 18 families with children and analyzed their conversation history with Google Home. They found that children viewed the smart assistant as a family member, more so than the adults did, and perceived the AI as capable of having thoughts and feeling emotions. Children between the ages of 5–7 developed emotional attachment to the device. In another study (Hiniker et al., 2021), researchers found that the way children spoke to a voice assistant could carry over into how they spoke to parents. In one example, a parent described, “If I do not listen to what my son is saying, he will just start shouting in an aggressive tone. He thinks, as Google responds to such a tone, I would too.” Other researchers (Beneteau et al., 2020) found that while these AI voice assistants could introduce conflict within families, they could also foster communication between family members, and adults used them to augment their parenting strategies.
@@ 1.4 Perceiving mind in chatbots
A growing body of work focuses on how people attribute the property of humanness to others, particularly AI (Dubosc et al., 2021; Y. Kim & Sundar, 2012; Seeger & Heinzl, 2018; Severson & Woodard, 2018; Waytz et al., 2010). Previous evidence suggests that attributing humanlike characteristics to artificial agents impacts how people interact with these agents (Blut et al., 2021; Damholdt et al., 2023; Guingrich & Graziano, 2024; Kühne & Peter, 2022; Pizzi et al., 2023; Rhim et al., 2022; Seeger & Heinzl, 2018; Stein et al., 2022). Viewing an AI agent as having a more sophisticated mind increases people’s trust in the agent and intent to use it, for example (Young & Monroe, 2019). Clifford Nass’s theory of human–computer interaction, the Computers Are Social Actors framework, first used the concept of ethopoeia to explain why people treat nonhuman agents as social actors (Nass et al., 1994). Ethopoeia involves attributing one’s own humanlike characteristics to nonhuman agents, which spurs the automatic use of human social scripts during interaction with such agents. Whether AI is or can ever be conscious is a contentious question (Chalmers, 2023; Graziano, 2023; Koch, 2019; Shanahan, 2024). Regardless, evidence indicates that at least some users perceive consciousness and other humanlike attributes in AI (Colombatto & Fleming, 2023), especially with AI that looks and acts humanlike (Broadbent et al., 2013; Marchesi et al., 2022; Stein & Ohler, 2017; Stein et al., 2020), and this perception may influence whether AI has a positive or negative psychological impact on people. In consciousness and Theory of Mind literature, some researchers suggest that consciousness is inherently an attribution, and that attributing similar consciousness to oneself and to others can facilitate understanding and foster relationships between people (Frith, 2002; Graziano, 2013; Prinz, 2017).
@@ 1.5 The current study
The goal of our study was to gain better insight into why people have relationships with chatbots and to understand the social consequences of human–chatbot relationships. We also aimed to address how perceptions of human likeness and mind in the chatbot played a role in social outcomes. We used an online survey to study the responses of people who were regular users of companion chatbots and those who were not. Sampling these populations allowed us to contrast an AI-user sample to a general one and to uncover possible cross-group similarities. We hypothesized that people’s opinions about companion chatbots would be significantly related to their perceptions of whether the chatbots had consciousness, agency, subjective experience, and overall human likeness. One possibility is that as people perceive chatbots to be more conscious and more humanlike, they may also rate them as more harmful, disturbing, and dangerous. This assumption reflects large-scale fears about AI advancements such as the singularity and p(doom) (Roose, 2023). Further, this negativity might stem from the fear that AI will infiltrate traditionally human roles or threaten human uniqueness (Ferrari et al., 2016). It might also stem from an uncanny valley effect, which was first introduced by Mori (2012) in 1970 and has since been a focus of human–AI interaction research. As an AI agent’s human likeness increases up to a certain threshold, people tend to view the agent more favorably. Once an AI agent’s appearance, mental capacities, or other human capabilities reach a certain threshold of too human but not human enough, people report feeling eerie or uneasy (Ciechanowski et al., 2019; Kätsyri et al., 2015; der Pütten & Krämer, 2014; der Pütten et al., 2019; Schein & Gray, 2015; Stein & Ohler, 2017). A second possibility, however, is that as people perceive chatbots to be more conscious and more humanlike, they may rate them as more beneficial because this allows them to engage in more fruitful social and emotional interaction.
@@ 2 Methods
Full ethical approval for the study and procedures was granted by Princeton University’s Internal Review Board. All participants provided informed consent and were paid \$4.00. We conducted the study online and sampled two groups: regular users of a companion chatbot (N=82: 69.5% men, N=57; 22% women, N=18; 2.4% nonbinary/other, N=2; 6.1% prefer not to say, N=5; age range 18–65+) and US and UK residents who were not companion chatbot users (N=135: 47.4% women, N=64; 42.2% men, N=57; 1.5% nonbinary/other, N=2; 8.9% prefer not to say, N=12; age range 18–65+). The majority of participants in both samples resided in the US and the UK (user sample: US=65.9%, other=29.3%, undisclosed=4.9%; non-user sample: UK=60%, US=32.6%, undisclosed=7.4%). We collected our data between January and February 2023.
@@
The survey given to the group of companion chatbot users contained 31 multiple choice and three free response questions. For the 31 multiple choice questions, the choices took the form of a 1-to-7 Likert scale. For the online survey given to participants, questions 1–21 appeared on page one of the survey; questions 22–31 appeared on page two; and three free-response questions appeared on page three. Five groups of questions were designed to assess specific psychological constructs that subjects may have applied to the companion chatbot. The scores for questions 3–5 were averaged together to provide an overall assessment of the subjects’ social health (“social health” index). We operationalized social health impacts as participants’ perceptions of how the chatbot relationship impacted their general social interactions, their relationships with family and friends, and their self- esteem. The scores for questions 6–11 were averaged together to provide an assessment of the subjects’ perception of how much the companion chatbot has a subjective experience (“experience” index). The scores for questions 12–15 were averaged together to provide an assessment of whether subjects believed the companion chatbot had a conscious mind (“consciousness” index). The scores for questions 17–21 were averaged together to provide an assessment of how much subjects considered the companion chatbot to have the properties of an active agent (“agency” index). Finally, the scores for questions 22–28 were averaged together to provide an assessment of how humanlike subjects believed the companion chatbot to be (“human-likeness” index). These groupings of questions were based on modifications of prior studies assessing perceptions of social health and consciousness, as well as preexisting indices for measuring human likeness and Theory of Mind ascriptions (Bartneck et al., 2009; K. Gray et al., 2011; Graziano, 2013; Ward et al., 2013).
The non-user group received a survey that was the same as the companion chatbot user group’s except in two ways. First, the survey for the non-user group began with an explanatory paragraph about Replika with images of the interface. Second, all questions were phrased as hypotheticals for non-users, who were asked to imagine themselves as users of the chatbot app. For example, the companion chatbot user group was asked, “Please rate how harmful or helpful your relationship with your Replika has been for your…” whereas the non-user group was asked, “Please rate how harmful or helpful you think your relationship with your Replika would be for your 1. Social interactions, 2. Relationships with family or friends, and 3. Self-esteem.”
All survey materials, anonymized data, and code used for data analysis are publicly available on our project’s OSF page (link).